OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence? A. 3, 7, 11 B. 4, 8, 12 C. 3, 6, 9 D. 4 , 7, 11

OpenStudy (anonymous):

4(1)-1=? 4(2)-1=? 4(3)-1=? haha easier?

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yup :D thanks :)